OpenStudy (anonymous):

convert the decimal to a fraction0.55. is the answer 11/20

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

thx just dbl. checking... lol

OpenStudy (anonymous):

easy way to check is to plug it into a calculator

OpenStudy (anonymous):

how do i sh.wk.

OpenStudy (anonymous):

.55 = 55/100 = 11/20

OpenStudy (anonymous):

ok. i did something wrong when i attempted it,thx =)